# Continual Lie algebras determined by chain complexes

### v5

@@ 2.2. The algebra of graded differential relations
for the sequence starting from $(5)$:
$n=n_{0}+n_{j}^{(LLK_{j+2})}-r_{j}^{(LLK_{j+2})},$ (2.16)
$i$, $j\geq 2$.
@@
One can easily see that not all elements in (2.2) are independent. For instance, from (2.8) and (2.11) we obtain $\left(n^{(L)}_{1}-r^{(L)}_{1}\right)=-\left(n_{1}^{(R)}-r^{(R)}_{1}\right)=n-n_{0}-1$. From (2.7)-(2.10), and from $(3^{\prime})$-$(3^{\prime\prime})$ of (2.2) we infer that $\alpha_{1}^{(L)}$, $\alpha_{1}^{(R)}$, and $\alpha_{i}^{(LLK_{i})}$, $\alpha_{i}^{(RRK_{i})}$ are related by a conjugation with respect to the product (2.2): $\Phi=\alpha_{1}^{(L)}\cdot\left(\Phi\cdot\alpha_{1}^{(R)}\right)$, $\delta(n_{0})\chi=\left(\alpha_{i}^{(LLK_{i})}\cdot\delta(n_{0})\;\chi\right)\cdot\alpha_{i}^{(RRK_{i})}$. Similar relations apply among other elements $\alpha_{i}^{(K_{i})}$. The sequence of relations (2.2) cancels when one of relations (2.8)–(2.11) or (2.15)–(2.16) for a sequence of pairs of equations is not fulfilled. The natural grading is given by the condition that both sides of differential relations in (2.2) belong to the same chain complex space. In this paper we consider the simplest form (2.1) of a complex. In general, for more complicated actions of $\delta$, such that $\delta(i):C(i)\to C(i+k(i))$, where $k(i)$ depends on $i\in\mathbb{Z}$ (see, e.g., [11] for non-trivial actions of certain $\delta$ operators among chain complex spaces for vertex algebras. Consideration of such complexes will be given by the author in another article). The corresponding differential relations as well as compatibility relations could be different from (2.2)
@@
As an upshot, the orthogonality condition (2.5) for all choices of $n_{0}$, $n\in\mathbb{Z}$, and the conditions (2.15)–(2.16), applied to the chain complex (2.1) bring about the structure of a graded algebra with differential relations (2.2) with respect to the multiplication (2.4). As we can see, the system of relations (2.2) has a tree structure. ”Left” and ”right” directions have a mixture of dependent elements. Let us denote $n(i)=n_{i}^{(K_{i})}$. For $n_{0}$, $n\in\mathbb{Z}$, let $I(n_{0},n)$ be the set of sequences of indices $(n(i))$, $i\geq 0$, marking all paths $(K_{i})$ in the tree structure of (2.2), describing ”left” or ”right” choice at each point. Let $I=\bigcup\limits_{n_{0},n\in\mathbb{Z}}\;I(n_{0},n)$, be the space of all sequences over the tree graph for a complex (2.1). Denote by $I_{0}\subset I$ the subset of such paths that include pairs of independent elements only. Then we are able to single out generators and commutation relations of a Lie algebra with a continual space of roots, (see subsection (1.1) and [17]).
@@ 2.3. Construction of a continual Lie algebra from a chain complex
Let us further assume that the spaces $C(i)$, $i\in\mathbb{Z}$ admit also an ordinary (not necessary commutative) product $\Phi\Psi$ among elements for $\Phi_{k}\in C(k)$, and $\Phi_{l}\in C(l)$, for all $k$, $l\in\mathbb{Z}$. Then, as a product (2.2), satisfying conditions (2.4) and (2.3), one can take
$\Phi_{k}\cdot\Phi_{l}=\left[\Phi_{k},\Phi_{l}\right]=\Phi_{k}\Phi_{l}-\Phi_{l}\Phi_{k},$ (2.17)
where brackets mean the ordinary commutator. It is known that the introduction of the commutator with respect to the original multiplication of an algebra transfers it into a Lie algebra when the Jacobi conditions are satisfied. In our case we show that the differential algebra $\mathcal{C}$ defined above being supplied with the orthogonality conditions deliver the structure of a continual Lie algebra.
The setup of this subsection combined with the previous subsection provides us with a proof of the main result of this paper, Theorem 1.
Though the structure of the system (2.2) may seem to be not very complicated, actual properties of the corresponding continual Lie algebra depends on properties of the spaces of a specific the bicomplex (2.1) and the nature of parameters $\Theta_{i}$. For a fixed choice of a path of independent functions/differential equations in the system (2.2), there exists a variety of choices on how to identify generators of a continual Lie algebra with generators of the differential algebra. Therefore, the actual form of commutation relations for the corresponding continual Lie algebra varies accordingly. One can also chose various ways how to define a grading for each specific $\mathcal{G}(\Theta_{n(i)})$, for generators of a continual Lie algebra (see subsection 1.1) resulting from (2.2). The structure of the product (2.2), the condition (2.3), and the action of the differentials $\delta$ provide the Jacobi identity for generators on the continual Lie algebra $\mathcal{G}(\Theta_{n(i)})$, and, simultaneously, apply conditions of the form (1.1) to elements of the parameter spaces $\Theta_{(n(i))}$, $i\geq 0$. In the next Section we specify the above construction and Theorem (1) in the case of double cochain complex (3.1)–(3.2) associated with differential forms [5]. We derive explicitly the generators and commutation relations for the corresponding continual Lie algebras.
@@ 3. An example: double complex associated with foliations
$C^{p,q}(\mathcal{F})=\prod_{U_{0}\stackrel{{\scriptstyle h_{1}}}{{\longrightarrow}}\cdots\stackrel{{\scriptstyle h_{p}}}{{\longrightarrow}}U_{p}}\Omega^{q}(U_{0}),$ (3.1)
where the product ranges over all $p$-tuples of holonomy embeddings $h_{i}$, $0\leq i\leq p$, between transversal sections from a fixed transversal basis $\mathcal{U}$, and $\Omega^{q}$ is the space of differential forms of order $q$. The vertical differential is defined as $(-1)^{p}d:C^{p,q}(\mathcal{F})\to C^{p,q+1}(\mathcal{F})$, where $d$ is the usual de Rham differential. The horizontal differential $\delta:C^{p,q}(\mathcal{F})\to C^{p+1,q}(\mathcal{F})$, is given by $\delta=\sum\limits_{i=0}^{k+1}(-1)^{i}\delta_{i}$, where
$\delta_{i}\omega(h_{1},\ldots,h_{k+1})=\left\{\begin{array}[]{lll}h_{1}^{*}\omega(h_{2},\ldots,h_{k+1}),\;\mbox{if $i=0,$}\\ \omega(h_{1},\ldots,h_{i+1}h_{i},\ldots,h_{k+1}),\;\mbox{if $0<i<k+1,$}\\ \omega(h_{1},\ldots,h_{k}),\;\mbox{if $i=k+1.$}\end{array}\right.$ (3.2)
@@
for $\omega\in C^{n,m}$ and $\eta\in C^{n^{\prime},m^{\prime}}$, and $h^{*}_{i}$ being the dual to $h_{i}$. Thus $(\omega\;\eta)(h_{1},\ldots,h_{n+n^{\prime}})\in C^{n+n^{\prime},m+m^{\prime}}$, and the product (3.3) delivers the structure of a bigraded differential algebra. The cohomology of this complex is called the Čech-de Rham cohomology ${\check{H}}^{*}_{\mathcal{U}}(M/\mathcal{F})$ of the leaf space $M/\mathcal{F}$ with respect to the transversal basis $\mathcal{U}$.
@@ 3.2. The double complex
In the previous subsection we recalled the formulation of the Čech-de Rham cohomology given in [5] for foliations on smooth manifolds (see subsection 3.1). In this case the spaces in (2.1) are differential forms $C^{n,m}(\mathcal{F})$ defined on a foliation, and the coboundary operators is given by $\delta^{p,q}=(-1)^{p}d+\delta^{q}$, where $d$ and $\delta$ operators are defined in subsection 3.1. The ordinary product for differential forms $\omega^{n,m}(h_{1},\ldots,h_{m})\in C^{n,m}\left(\mathcal{F}\right)$ is given by (3.3). The product (2.2) required for the formulation of Section 2 is provided by the commutator (2.17). As it was mentioned in [5], the bicomplex (3.1)–(3.2) has the structure of a bigraded differential algebra with respect to the ordinary product (3.3). According to explanations of Section 2, for the bicomplex (3.1)–(3.2) generate the graded differential algebra with relations (2.2) with respect to the product (2.17).
Let us recall that we assume non-negative indices for all bicomplex spaces $C^{n_{i},m_{i}}$. Let $\chi\in C^{n_{0},m_{0}}$, $\Phi\in C^{n,m}$. The orthogonality condition (2.5) with respect ot the product (2.17) leads to systems of the form (2.2) when applied to the double complex (2.1). In particular, for $\chi\in C^{n_{0},m_{0}}$ and $\Phi\in C^{n_{0},m_{0}}$, we obtain the system (2.2) of relations for elements $\chi$, $\delta^{n_{0},m_{0}}\chi$, $\Phi$, $\delta^{n,m}\Phi$, $\alpha_{i}^{(K_{i})}\in C^{n_{i}^{(K_{i})},m_{i}^{(K_{i})}}$, $\delta^{n(i),m(i)}\alpha_{i}^{(K_{i})}$, $i\geq 1$, for $n\geq 0$, $m\geq 0$. Let $r^{(K_{i})}_{i}$ and $t^{(K_{i})}_{i}$ be numbers of common degrees and transversal sections of $C^{n_{i},m_{i}}$ for the forms $\chi\in C^{n_{0},m_{0}}$ and $\Phi\in C^{n_{0},m_{0}}$. Then the compatibility conditions (2.8)–(2.11) and (2.15)–(2.16) for indices $n$, $m$, $n_{0}$, $m_{0}$, $n_{i}^{(K_{i})}$, $m_{i}^{(K_{i})}$, satisfy the relations in vector form:
@@
$i$, $j\geq 2$. Note that it is assumed that all the resulting indices in the compatibility conditions are non-negative. For the general complex (2.1), $i\in\mathbb{Z}$. For the bicomplex (3.1)–(3.2) we have $p$, $q$ are non-negative. Since we assume that $0\leq r^{(LLK_{i})}_{i}\leq n^{(LLK_{i})}_{i}$, and $0\leq t^{(LLK_{i})}\leq m^{(LLK_{i})}_{i}$, from the compatibility conditions (3.4)–(3.5) and (3.6)–(3.7) we see that, depending on the signs of $n-n_{0}-\delta_{i,1}$, $m-m_{0}-\delta_{i,1}$, and only one branch of systems of the form $(2)$$(2^{\prime})$ exists at each vertex of the three graph associated to the double complex (3.1)–(3.2).
@@
As explained at the end of Section 2, due to Theorem (1), a path $(K_{i})$, $i\geq 1$, defining the generators $\left\{\chi,\;\delta^{n_{0},m_{0}}\chi,\;\Phi,\;\delta^{n,m}\Phi,\;\alpha_{i}^{(K_{i})},\;\delta^{n(i),m(i)}\alpha_{i}^{(K_{i})}\right\}$, $n(i)=n_{i}^{(K_{i})}$, $m(i)=m_{i}^{(K_{j})}$, and relations for independent elements of $C^{n,m}$ in (2.2), form a continual Lie algebra $\mathcal{G}(\mathcal{F})$ with the space of roots provided by the holonomy mappings and of parameters of the foliation $\mathcal{F}$.

### v6

@@ 2.2. The algebra of graded differential relations
for the sequence starting from $(5)$ for $i$, $j\geq 2$
$n=n_{0}+n_{j}^{(LLK_{j+2})}-r_{j}^{(LLK_{j+2})}.$ (2.16)
One can easily see that not all elements in (2.2) are independent. For instance, from (LABEL:condi1) and (LABEL:condi2) we obtain $\left(n^{(L)}_{1}-r^{(L)}_{1}\right)=-\left(n_{1}^{(R)}-r^{(R)}_{1}\right)=n-n_{0}-1$. From (LABEL:trojka)-(2.10), and from $(3^{\prime})$-$(3^{\prime\prime})$ of (2.2) we infer that $\alpha_{1}^{(L)}$, $\alpha_{1}^{(R)}$, and $\alpha_{i}^{(LLK_{i})}$, $\alpha_{i}^{(RRK_{i})}$ are related by a conjugation with respect to the product (LABEL:initialproduct): $\Phi=\alpha_{1}^{(L)}\cdot\left(\Phi\cdot\alpha_{1}^{(R)}\right)$, $\delta(n_{0})\chi=\left(\alpha_{i}^{(LLK_{i})}\cdot\delta(n_{0})\;\chi\right)\cdot\alpha_{i}^{(RRK_{i})}$. Similar relations apply among other elements $\alpha_{i}^{(K_{i})}$.
We supply here a recursive representation summarizing the systems (2.2)-(2.2) in the following form. In order to establish a clear way to rewrite such a system we make the recursion explicit over the word $K_{i}$. The indexing set consists of the following. Let $\mathcal{K}=\{L,R\}^{*}$ be the set of all finite words over $\{L,R\}$, and let $\epsilon$ denote the empty word. The generators are the following. For each $i\geq 1$, define
$\alpha_{i}:\mathcal{K}\to\bigsqcup_{n\in\mathbb{Z}}C(n),\qquad n_{i}:\mathcal{K}\to\mathbb{Z},\qquad r_{i}:\mathcal{K}\to\mathbb{Z}_{\geq 0}.$
Now we proceed with levels of the system (2.2)-(2.2). The first level ($i=1$) is determined by (LABEL:trojka) and (2.10):
$\delta(n_{0})\chi=\Phi\cdot\alpha_{1}^{(R)},\qquad\Phi=\alpha_{1}^{(L)}\cdot\delta(n_{0})\chi.$
The indices satisfy the compatibility conditions:
$n_{0}+1=n+n_{1}^{(R)}-r_{1}^{(R)},\qquad n=n_{1}^{(L)}+n_{0}+1-r_{1}^{(L)}.$
The recursion: let $K\in\mathcal{K}$. Then we have the R-extension: if $\alpha_{i}^{(K)}$ appears in a relation of the form
$\delta(n_{0})\chi=(\cdots)\cdot\alpha_{i}^{(K)},$
then applying $\delta(\cdots)$ and using the graded Leibniz rule (LABEL:leibniz0) gives
$0=(\cdots)^{\prime}\cdot\alpha_{i}^{(K)}+(-1)^{\deg(\cdots)\deg(\alpha_{i}^{(K)})}(\cdots)\cdot\delta\bigl(n_{i}^{(K)}\bigr)\alpha_{i}^{(K)}.$
Hence there exists $\alpha_{i+1}^{(KR)}$ such that
$\delta(n_{0})\chi=(\cdots)^{\prime}\cdot\alpha_{i+1}^{(KR)}.$
Thus $\alpha_{i+1}^{(KR)}$ is defined by $\delta(n_{0})\chi=(\delta(\cdots))\cdot\alpha_{i+1}^{(KR)}$. The indices satisfy (LABEL:rel2)
$n_{0}=n+n_{i+1}^{(KR)}-r_{i+1}^{(KR)}.$
We have also the L-extension. If $\alpha_{i}^{(K)}$ appears in a relation of the form
$\delta(n)\Phi=\alpha_{i}^{(K)}\cdot\delta(n_{0})\chi,$
then applying $\delta(\cdots)$ gives
$0=\delta\bigl(n_{i}^{(K)}\bigr)\alpha_{i}^{(K)}\cdot\delta(n_{0})\chi.$
Hence there exists $\alpha_{i+1}^{(KL)}$ such that
$\delta(n)\Phi=\alpha_{i+1}^{(KL)}\cdot\delta(n_{0})\chi.$
Thus, $\alpha_{i+1}^{(KL)}$ is defined by $\delta(n)\Phi=\alpha_{i+1}^{(KL)}\cdot\delta(n_{0})\chi$. The indices satisfy (LABEL:rel3)
$n=n_{0}+n_{i+1}^{(KL)}-r_{i+1}^{(KL)}.$
The explicit recursive formulation: for all $K\in\mathcal{K}$, $\alpha_{i+1}^{(KR)}$ is the unique element such that
$\delta(n_{0})\chi=(\delta^{(\cdots)}\Phi)\cdot\alpha_{i+1}^{(KR)},$
$\alpha_{i+1}^{(KL)}$ is the unique element such that
$\delta(n)\Phi=\alpha_{i+1}^{(KL)}\cdot\delta(n_{0})\chi.$
The compact operator form of the system is the following: define maps $T_{R}(\alpha)=\alpha^{\prime}$ such that
$\delta(n_{0})\chi=(\delta\Phi)\cdot\alpha^{\prime},$
and $T_{L}(\alpha)=\alpha^{\prime}$ such that
$\delta(n)\Phi=\alpha^{\prime}\cdot\delta(n_{0})\chi.$
Then for $K=k_{1}\cdots k_{m}$,
$\alpha_{i}^{(K)}=T_{k_{m}}\circ\cdots\circ T_{k_{1}}(\alpha_{1}^{(\cdot)}).$
All that can be summarized in the table:
$\begin{array}[]{|c|c|c|c|}\hline\cr\textbf{Level}&\textbf{Branch}&\textbf{Relation}&\textbf{Recursive form}\\ \hline\cr 0&-&0=\Phi\cdot\delta(n_{0})\chi&\mathcal{R}_{0}=\Phi\perp\delta(n_{0})\chi\\ \hline\cr 1&R&\delta(n_{0})\chi=\Phi\cdot\alpha^{(R)}_{1}&\alpha^{(R)}_{1}\in C(n^{(R)}_{1})\\ 1&L&\Phi=\alpha^{(L)}_{1}\cdot\delta(n_{0})\chi&\alpha^{(L)}_{1}\in C(n^{(L)}_{1})\\ \hline\cr 2&RR&\delta(n_{0})\chi=(\delta(n)\Phi)\cdot\alpha^{(RR)}_{2}&\alpha^{(RR)}_{2}=\mathcal{F}_{R}(\alpha^{(R)}_{1})\\ 2&LL&\delta(n)\Phi=\alpha^{(LL)}_{2}\cdot\delta(n_{0})\chi&\alpha^{(LL)}_{2}=\mathcal{F}_{L}(\alpha^{(L)}_{1})\\ \hline\cr i\geq 1&RK_{i}&\delta(\cdot)=(\delta(\cdot))\cdot\alpha^{(RK_{i})}_{i+1}&\alpha^{(RK_{i})}_{i+1}=\mathcal{F}_{R}\big(\alpha^{(K_{i})}_{i}\big)\\ i\geq 1&LK_{i}&\delta(\cdot)=\alpha^{(LK_{i})}_{i+1}\cdot\delta(\cdot)&\alpha^{(LK_{i})}_{i+1}=\mathcal{F}_{L}\big(\alpha^{(K_{i})}_{i}\big)\\ \hline\cr\end{array}$
$\alpha^{(K_{i+1})}_{i+1}=\begin{cases}\mathcal{F}_{R}\big(\alpha^{(K_{i})}_{i}\big),&K_{i+1}=(K_{i},R),\\[6.0pt] \mathcal{F}_{L}\big(\alpha^{(K_{i})}_{i}\big),&K_{i+1}=(K_{i},L),\end{cases}$
with induced relations: the right branch: $\delta(\cdot)=(\delta(\cdot))\cdot\alpha^{(K_{i})}_{i}$, the left branch: $\delta(\cdot)=\alpha^{(K_{i})}_{i}\cdot\delta(\cdot)$. The system forms a binary tree of relations. Each node corresponds to an element $\alpha_{i}^{(K_{i})}$. Paths $K_{i}=(L,R)$ encode iterated left/right factorizations. (2.13) (2.2) corresponds to the right subtree , $RRK_{i}$ (2.14) (2.2) to the left subtree $LLK_{i}$.
@@
The sequence of relations (2.2) cancels when one of relations (LABEL:condi1)-(LABEL:condi2) or (LABEL:rel2)-(LABEL:rel3) for a sequence of pairs of equations is not fulfilled. The natural grading is given by the condition that both sides of relations in (2.2) belong to the same chain complex space. In this paper we consider the simplest form (LABEL:complex) of a complex. In general, for more complicated actions of $\delta$, such that $\delta(i):C(i)\to C(i+k(i))$, where $k(i)$ depends on $i\in\mathbb{Z}$ (see, e.g., [11] for non-trivial actions of certain $\delta$ operators among chain complex spaces for vertex algebras. Consideration of such complexes will be given by the author in another article). The corresponding relations as well as compatibility relations could be different from (2.2)
@@
As an upshot, the orthogonality condition (LABEL:ortho) for all choices of $_{0}, nn\in\mathbb{Z}$, and the conditions (LABEL:rel2)-(LABEL:rel3), applied to the chain complex (LABEL:complex) bring about the structure of a graded algebra with relations (2.2) with respect to the multiplication (LABEL:prodcond). As we can see, the system of relations (2.2) has a tree structure. ”Left” and ”right” directions have a mixture of dependent elements. Let us denote $n(i)=n_{i}^{(K_{i})}$. For $_{0}, nn\in\mathbb{Z}$, let $I(n_{0},n)$ be the set of sequences of indices $(n(i))$, $i\geq 0$, marking all paths $(K_{i})$ in the tree structure of (2.2), describing ”left” or ”right” choice at each point. Let $I=\bigcup\limits_{n_{0},n\in\mathbb{Z}}\;I(n_{0},n)$, be the space of all sequences over the tree graph for a complex (LABEL:complex). Denote by $I_{0}\subset I$ the subset of such paths that include pairs of independent elements only. Then we are able to single out generators and commutation relations of a Lie algebra with a continual space of roots, (see subsection (LABEL:continual) and [17]).
@@ 2.3. Construction of a continual Lie algebra from a chain complex
Let us further assume that the spaces $C(i)$, $i\in\mathbb{Z}$ admit an ordinary (not necessary commutative) product $\Phi\Psi$ among elements for $\Phi_{k}\in C(k)$, and $\Phi_{l}\in C(l)$, for all $k$, $l\in\mathbb{Z}$. Then, as a product (LABEL:initialproduct), satisfying conditions (LABEL:leibniz0) and (LABEL:prodcond), one can take
$\Phi_{k}\cdot\Phi_{l}=\left[\Phi_{k},\Phi_{l}\right]_{gr}=\Phi_{k}\Phi_{l}-(-1)^{\deg\left(\Phi_{k}\cdot\Phi_{l}\right)}\Phi_{l}\Phi_{k},$ (2.17)
where brackets mean the ordinary graded commutator. It is known that the introduction of the commutator with respect to the original multiplication of an algebra transfers it into a Lie algebra when the Jacobi conditions are satisfied. In our case we show that the differential algebra $\mathcal{C}$ defined above being supplied with the orthogonality conditions deliver the structure of a continual Lie algebra. The setup of this subsection combined with the previous subsection provides us with a proof of the main result of this paper, Theorem ??.
Though the structure of the system (2.2) may seem to be not very complicated, actual properties of the corresponding continual Lie algebra depend on properties of the spaces of a specific the bicomplex (LABEL:complex) and the nature of parameters $\Theta_{i}$. For a fixed choice of a path of independent functions/differential relations in the system (2.2), there exists a variety of choices on how to identify generators of a continual Lie algebra with generators of the differential algebra. Therefore, the actual form of commutation relations for the corresponding continual Lie algebra varies accordingly. One can also chose various ways how to define a grading for each specific $\mathcal{G}(\Theta_{n(i)})$, for generators of a continual Lie algebra (see subsection ??) resulting from (2.2). The structure of the product (LABEL:initialproduct), the condition (LABEL:leibniz0), and the action of the differentials $\delta$ provide the Jacobi identity for generators on the continual Lie algebra $\mathcal{G}(\Theta_{n(i)})$, and, simultaneously, apply conditions of the form (1.1) to elements of the parameter spaces $\Theta_{(n(i))}$, $i\geq 0$. In the next Section we specify the above construction and Theorem (LABEL:mainpro) in the case of double cochain complex (LABEL:Cpq)-(LABEL:deltas) associated with differential forms [5]. We derive explicitly the generators and commutation relations for the corresponding continual Lie algebras.
@@ 3. An example: double complex associated with foliations
$C^{p,q}(\mathcal{F})=\prod_{U_{0}\mathrel{{\mathop{\longrightarrow}\limits^{h_{1}}}}\cdots\mathrel{{\mathop{\longrightarrow}\limits^{h_{p}}}}U_{p}}\Omega^{q}(U_{0}),$ (3.1)
where the product ranges over all $p$-tuples of holonomy embeddings $h_{i}$, $0\leq i\leq p$, between transversal sections from a fixed transversal basis $\mathcal{U}$, and $\Omega^{q}$ is the space of differential forms of order $q$. The vertical differential is defined as $(-1)^{p}d:C^{p,q}(\mathcal{F})\to C^{p,q+1}(\mathcal{F})$, where $d$ is the usual de Rham differential. The horizontal differential $\mathcal{D}:C^{p,q}(\mathcal{F})\to C^{p+1,q}(\mathcal{F})$, is given by $\mathcal{D}=\sum\limits_{i=0}^{k+1}(-1)^{i}\mathcal{D}_{i}$, where
$\mathcal{D}_{i}\omega(h_{1},\ldots,h_{k+1})=\left\{\begin{array}[]{lll}h_{1}^{*}\omega(h_{2},\ldots,h_{k+1}),\;\mbox{if $i=0,$}\\ \omega(h_{1},\ldots,h_{i+1}h_{i},\ldots,h_{k+1}),\;\mbox{if $0<i<k+1,$}\\ \omega(h_{1},\ldots,h_{k}),\;\mbox{if $i=k+1.$}\end{array}\right$. (3.2)
@@
for $\omega\in C^{n,m}$ and $\eta\in C^{n^{\prime},m^{\prime}}$, and $h^{*}_{i}$ being the dual to $h_{i}$. Thus $(\omega\;\eta)(h_{1},\ldots,h_{n+n^{\prime}})\in C^{n+n^{\prime},m+m^{\prime}}$, and the product (LABEL:bigradif) delivers the structure of a bigraded differential algebra. The cohomology of this complex is called the Čech-de Rham cohomology ${\check{H}}^{*}_{\mathcal{U}}(M/\mathcal{F})$ of the leaf space $M/\mathcal{F}$ with respect to the transversal basis $\mathcal{U}$.
@@ 3.2. The double complex
In the previous subsection we recalled the formulation of the Čech-de Rham cohomology given in [5] for foliations on smooth manifolds (see subsection ??). In this case the spaces in (LABEL:complex) are differential forms $C^{n,m}(\mathcal{F})$ defined on a foliation, and the coboundary operators is given by $\delta^{p,q}=(-1)^{p}d+\delta^{q}$, where $d$ and $\delta$ operators are defined in subsection ??. The ordinary product for differential forms $\omega^{n,m}(h_{1},\ldots,h_{m})\in C^{n,m}\left(\mathcal{F}\right)$ is given by (LABEL:bigradif). The product (LABEL:initialproduct) required for the formulation of Section ?? is provided by the commutator (2.17). As it was mentioned in [5], the bicomplex (LABEL:Cpq)-(LABEL:deltas) has the structure of a bigraded differential algebra with respect to the ordinary product (LABEL:bigradif). According to explanations of Section ??, for the bicomplex (LABEL:Cpq)-(LABEL:deltas) generates a graded differential algebra with relations (2.2) with respect to the product (2.17).
Let us recall that we assume non-negative indices for all bicomplex spaces $C^{n_{i},m_{i}}$. Let $\chi\in C^{n_{0},m_{0}}$, $\Phi\in C^{n,m}$. The orthogonality condition (LABEL:ortho) with respect ot the product (2.17) leads to systems of the form (2.2) when applied to the double complex (LABEL:complex). In particular, for $\chi\in C^{n_{0},m_{0}}$ and $\Phi\in C^{n_{0},m_{0}}$, we obtain the system (2.2) of relations for elements $\chi$, $\delta^{n_{0},m_{0}}\chi$, $\Phi$, $\delta^{n,m}\Phi$, $\alpha_{i}^{(K_{i})}\in C^{n_{i}^{(K_{i})},m_{i}^{(K_{i})}}$, $\delta^{n(i),m(i)}\alpha_{i}^{(K_{i})}$, $i\geq 1$, for $n\geq 0$, $m\geq 0$. Let $r^{(K_{i})}_{i}$ and $t^{(K_{i})}_{i}$ be numbers of common degrees and transversal sections of $C^{n_{i},m_{i}}$ for the forms $\chi\in C^{n_{0},m_{0}}$ and $\Phi\in C^{n_{0},m_{0}}$. Then the compatibility conditions (LABEL:condi1)-(LABEL:condi2) and (LABEL:rel2)-(LABEL:rel3) for indices $n$, $m$, $n_{0}$, $m_{0}$, $n_{i}^{(K_{i})}$, $m_{i}^{(K_{i})}$, satisfy the relations in vector form:
@@
$i$, $j\geq 2$. Note that it is assumed that all the resulting indices in the compatibility conditions are non-negative. For the general complex (LABEL:complex), $i\in\mathbb{Z}$. For the bicomplex (LABEL:Cpq)-(LABEL:deltas) we have $p$, $q$ are non-negative. Since we assume that $0\leq r^{(LLK_{i})}_{i}\leq n^{(LLK_{i})}_{i}$, and $0\leq t^{(LLK_{i})}\leq m^{(LLK_{i})}_{i}$, from the compatibility conditions (3.4)-(3.5) and (3.6)-(3.7) we see that, depending on the signs of $n-n_{0}-\delta_{i,1}$, $m-m_{0}-\delta_{i,1}$, and only one branch of systems of the form $(2)(2^{\prime})$ exists at each vertex of the three graph associated to the double complex (LABEL:Cpq)-(LABEL:deltas).
@@
As explained at the end of Section ??, due to Theorem (LABEL:mainpro), a path $(K_{i})$, $i\geq 1$, defining the generators $\left\{\chi,\;\delta^{n_{0},m_{0}}\chi,\;\Phi,\;\delta^{n,m}\Phi,\;\alpha_{i}^{(K_{i})},\;\delta^{n(i),m(i)}\alpha_{i}^{(K_{i})}\right\}$, $n(i)=n_{i}^{(K_{i})}$, $m(i)=m_{i}^{(K_{j})}$, and relations for independent elements of $C^{n,m}$ in (2.2), form a continual Lie algebra $\mathcal{G}(\mathcal{F})$ with the space of roots provided by the holonomy mappings and of parameters of the foliation $\mathcal{F}$.